CS 582 — Database Management System II— Spring 2007

Homework 6 Deadline: 4/23/2007 – 11:55pm

Develop a XML schema for use in the banking system with the following information:

1. The bank might have more than one customers and many branches.

2. A branch is identified by their location (or address) and an internal unique routing number (12 digit integer).

3. A customers might have more than one account (12 digit integer).

4. The bank stores SSN (9 digits integer), Name (string), and Address of its customers (and of course, the accounts).

5. Addresses always have three parts: Street name, Street number, and Zip Code which are string, integer, and 5 digit integer respectively.

6. Each account has a number, a balance (integer), and is associated with some customers of the bank.

7. The customers can withdraw only from his accounts but can deposit on any account (think about a gift given to you:-) Withdraw is limited to the current balance. The information on withdraws and deposits made by a customer needs to be kept as well.

Prepare the XML schema and put it in your public html directory and submit the URL as the answer for the homework.